OpenStudy (anonymous):

f(x)= 8/x+2, g(x)= 8/7x. what is (f of g)(x)

OpenStudy (anonymous):

fg(x)=8/((8/7)x)+2 fg(x)=(7/x)+2 I hope you're clear of what I wrote :P

OpenStudy (anonymous):

it's simply 7x+2

OpenStudy (anonymous):

i need the write answer because this is for my math test